# STGFormer: Spatio-Temporal GraphFormer for 3D Human Pose Estimation in Video

### v1

@@ IV-B Results and Analysis
Performance Comparison on Human3.6M. We conducted a comparative analysis of our method against state-of-the-art approaches on the Human3.6M dataset. Tables I and II present our performance in terms of P1 and P2 errors, respectively, utilizing CPN pretrain 2D coordinates as inputs. Overall, our method has established a new state-of-the-art in a 243-frame setting, outperforming transformer-based, GCN-based, and graph-augmented transformer methods. Benefiting from our strong learning capability for body information, our approach demonstrated improvements over identical transformer-based methods such as StridedFormer [22], MHFormer [21], and MixSTE [20], with gains of 3.4mm, 2.7mm, and 0.6mm on P1 error, and 3.5mm, 2.7mm, and 0.9mm on P2 error, respectively. When compared to UGCN [42], GLA-GCN [11], and STCFormer [25], our method showed enhancements of 5.3mm, 4.1mm, and 0.7mm on P1 error. It is noteworthy that in a series of actions, 10 categories achieved the best-recorded performance to date.
Table III demonstrates the performance of our method on P1 error when using ground-truth 2D poses as input. Compared to CPN 2D poses, ground-truth 2D poses eliminate noise interference, allowing each model to achieve peak performance with ground-truth inputs. Overall, the trend of P1 error changes remains consistent with our method when compared to other methods. Our method reaches state-of-the-art performance at 243 frames input, with a P1 error of 21.4mm, leading the currently most advanced model, MixSTE [20], by 0.2mm.
Performance Comparison on MPI-INF-3DHP. To validate the generalizability of our proposed framework, we tested it on the MPI-INF-3DHP dataset, known for its complex backgrounds. Given the shorter video sequences, we configured the number of input frames at 9, 27, and 81, respectively. Detailed performance comparisons are presented in Table IV. Consistent with the results observed on Human3.6M, the optimal performance is achieved when the frame count is set to 81, achieving PCK of 98.8, AUC of 84.1, and P1 error of 19.2mm. This performance surpasses the current state-of-the-art model by 0.1 in PCK and 0.2 in AUC. These findings highlight the efficacy of the proposed method in handling scenarios of increased complexity.
TABLE I: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN as inputs under Protocol 1, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
@@
StridedFormer [22]* (T=243) TMM’22 40.3 43.3 40.2 42.3 45.6 52.3 41.8 40.5 55.9 60.6 44.2 43.0 44.2 30.0 30.2 43.7
MHFormer [21]* (T=351) CVPR’22 39.2 43.1 40.1 40.9 44.9 51.2 40.6 41.3 53.5 60.3 43.7 41.1 43.8 29.8 30.6 43.0
MixSTE [20]* (T=243) CVPR’22 37.6 40.9 37.3 39.7 42.3 49.9 40.1 39.8 51.7 55.0 42.1 39.8 41.0 27.9 27.9 40.9
P-STMO [41]* (T=243) ECCV’22 38.9 42.7 40.4 41.1 45.6 49.7 40.9 39.9 55.5 59.4 44.9 42.2 42.7 29.4 29.4 42.8
HSTFormer [18]* (T=81) arXiv’23 39.5 42.0 39.9 40.8 44.4 50.9 40.9 41.3 54.7 58.8 43.6 40.7 43.4 30.1 30.4 42.7
Einfalt et al. [19]* (T=351) WACV’23 39.6 43.8 40.2 42.4 46.5 53.9 42.3 42.5 55.7 62.3 45.1 43.0 44.7 30.1 30.8 44.2
UGCN [42]^(T=96) ECCV’20 41.3 43.9 44.0 42.2 48.0 57.1 42.2 43.2 57.3 61.3 47.0 43.5 47.0 32.6 31.8 45.6
GLA-GCN [11]^(T=243) CVPR’23 41.3 44.3 40.8 41.8 45.9 54.1 42.1 41.5 57.8 62.9 45.0 42.8 45.9 29.4 29.9 44.4
STCFormer [25$\dagger$ (T=243) CVPR’23 39.6 41.6 37.4 38.8 43.1 51.1 39.1 39.7 51.4 57.4 41.8 38.5 40.7 27.1 28.6 41.0
STGFormer (Ours, T=81) 39.5 41.8 37.2 39.2 43.2 50.5 40.3 40.2 51.1 56.6 42.3 39.0 40.3 27.7 28.1 41.1
STGFormer (Ours, T=243) 38.4 40.7 37.1 38.1 42.7 49.9 39.0 39.3 50.4 56.0 41.4 38.3 39.1 26.9 28.1 40.3
TABLE II: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN as inputs under Protocol 2, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
@@
StridedFormer [22]* (T =243) TMM’22 32.7 35.5 32.5 35.4 35.9 41.6 33.0 31.9 45.1 50.1 36.3 33.5 35.1 23.9 25.0 35.2
MHFormer [21]* (T=351) CVPR’22 31.5 34.9 32.8 33.6 35.3 39.6 32.0 32.2 43.5 48.7 36.4 32.6 34.3 23.9 25.1 34.4
MixSTE [20]* (T=243) CVPR’22 30.8 33.1 30.3 31.8 33.1 39.1 31.1 30.5 42.5 44.5 34.0 30.8 32.7 22.1 22.9 32.6
P-STMO [41]* (T=243) ECCV’22 31.3 35.2 32.9 33.9 35.4 39.3 32.5 31.5 44.6 48.2 36.3 32.9 34.4 23.8 23.9 34.4
HSTFormer [18]* (T=81) arXiv’23 31.1 33.7 33.0 33.2 33.6 38.8 31.9 31.5 43.7 46.3 35.7 31.5 33.1 24.2 24.5 33.7
Einfalt et al. [19]* (T=351) WACV’23 32.7 36.1 33.4 36.0 36.1 42.0 33.3 33.1 45.4 50.7 37.0 34.1 35.9 24.4 25.4 35.7
UGCN [42]^(T=96) ECCV’20 32.9 35.2 35.6 34.4 36.4 42.7 31.2 32.5 45.6 50.2 37.3 32.8 36.3 26.0 23.9 35.5
GLA-GCN [11]^(T=243) CVPR’23 32.4 35.3 32.6 34.2 35.0 42.1 32.1 31.9 45.5 49.5 36.1 32.4 35.6 23.5 24.7 34.8
STCFormer [25$\dagger$ (T=243) CVPR’23 29.5 33.2 30.6 31.0 33.0 38.0 30.4 29.4 41.8 45.2 33.6 29.5 31.6 21.3 22.6 32.0
STGFormer (Ours, T=81) 29.6 33.2 30.2 31.3 33.0 37.7 30.1 30.4 41.7 44.6 33.5 29.5 31.7 22.0 23.1 32.1
STGFormer (Ours, T=243) 29.3 33.0 29.9 30.6 32.5 37.5 30.3 29.6 41.0 44.1 33.2 29.1 30.9 21.7 22.8 31.7
TABLE III: Performance Comparison on Human3.6M takes the ground-truth 2D pose as inputs under Protocol 1, where Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
@@
TABLE IV: Performance Comparison on MPI-INF-3DHP: Mark * denotes transformer-based methods, Mark ^indicates GCN-based methods, and Mark $\dagger$ represents graph-augmented transformer methods.
@@
MotionAGFormer [43$\dagger$ (T=27) WACV’23 98.2 83.5 19.2
@@
In this section, we conduct a qualitative analysis of the model’s performance by visualizing the attention layers and assessing the effects of model on Human3.6M dataset.
Attention visualization. We visualized the final spatial and temporal attention layers in our module, as shown in Figure 3. As expected, Figure 4(a) shows that the spatial attention layer map indicates that our proposed model has learned various patterns of different joints during movements. On the temporal attention layer, Figure 4(b) reveals a strong correlation between joints across adjacent frames and higher-order frames.
Figure 4: Visualization of attention layer maps from the spatial and temporal attention modules in the STGFormer.
Result visualization on Human3.6M. The visualization of our model’s performance on the Human3.6M dataset is depicted in Figure 5, where four examples from the Walking, Sitting, Greeting, and Eating sequences were randomly selected. Due to the framework’s strong modeling capabilities in both temporal and spatial dimensions for human action sequences, our method exhibits robust performance in complex actions such as Sitting (second row) and Eating (fourth row), effectively reconstructing poses in three-dimensional space.
@@
To further investigate the proposed model, we conducted a series of ablation studies on the Human3.6M dataset using CPN-pretrained 2D keypoint coordinates as input. The first set of ablation studies focused on how the performance of our proposed model varies under different frame counts. As illustrated in Table V, the P1 error comparison of our model with other methods on the Human3.6M dataset is detailed for 27, 81, and 243 frames. Overall, we observed a trend of decreasing P1 error with an increasing frame count in our method and others. Among these methods, our model achieved the best results at 27, 81, and 243 frames, demonstrating strong performance in 3D pose estimation in long sequence videos. Notably, with a comparable number of parameters, our model requires less than 79% of the FLOPs compared to the currently best-performing MixSTE [20] at 243 frames.
TABLE V: The P1 error comparisons with different numbers of frame on Human3.6M dataset. The best result in each column is highlighted in red.
@@
The second set of ablation studies examined the effects of various components within our proposed model on its performance, utilizing CPN-pretrained 2D coordinate data from 27 frames as input. Initially, spatial graph attention (SGA) and temporal graph attention (TGA) were analyzed separately, with MHR-GCN components replaced by an MLP. Subsequently, only the combined spatial-temporal graph attention was operational in the STGA-only set. Spatial MHR-GCN (SMHR-GCN) and Temporal MHR-GCN (TMHR-GCN) constitute the MHR-GCN components. Table VI offers detailed insights into how each component influences the overall performance of our model. Employing either spatial or temporal graph attention alone resulted in a substantial increase in P1 error, reaching 263.2mm and 57.8mm, respectively. By integrating both spatial and temporal graph attention, and adding SMHR-GCN and TMHR-GCN, the errors were further reduced by 6.4mm and 8.3mm, respectively.
TABLE VI: Performance contribution of each component in the proposed framework on Human3.6M dataset.
STGA SMHR-GCN TMHR-GCN P1 (mm)
SGA 263.2
@@
STGA $\checkmark$ 53.3
+SMHR-GCN $\checkmark$ $\checkmark$ 46.9
+TMHR-GCN $\checkmark$ $\checkmark$ 45.0
+MHR-GCN $\checkmark$ $\checkmark$ $\checkmark$ 43.7
V Conclusion
In this paper, we introduce the Spatio-Temporal GraphFormer, a novel model that integrates long-range temporal-spatial dependencies with graph structural information in its attention mechanisms and utilizes a dual-path modulated hop-wise regular GCN Module to enhance the extraction of graph information for 3D human pose estimation in videos. Specifically, the spatio-temporal criss-cross graph attention separately models spatial and temporal correlations, leveraging encoded graph structure information within the model. The dual-path modulated hop-wise regular GCN Module adjusts graph weights and accounts for high-order dependencies across both time and space dimensions. Experiments on two benchmarks affirm the efficacy and superior generalization capability of our proposed model over state-of-the-art techniques. For future work, we plan to develop a multi-person skeletal sequence spatio-temporal graph representation method to extend our model’s applicability to 3D human pose estimation in multi-person scenarios and broaden its use in various contexts.

### v2

@@ 4.2 Results and Analysis
Performance Comparison on Human3.6M. We conducted a comparative analysis of our method against state-of-the-art approaches on the Human3.6M dataset. Tables 1 and 2 present our performance in terms of P1 and P2 errors, respectively, using CPN and SH pre-trained 2D coordinates as inputs. Our method establishes a new state-of-the-art for a 243-frame setting, outperforming transformer-based, GCN-based, and graph-augmented transformer methods. For CPN, benefiting from our model’s strong ability to learn body information, our approach shows improvements over transformer-based methods such as StridedFormer [22], MHFormer [21], and MixSTE [20], with gains of 3.4 mm, 2.7 mm, and 0.6 mm in P1 error, and 3.5 mm, 2.7 mm, and 0.9 mm in P2 error, respectively. Compared to UGCN [42], GLA-GCN [11], and STCFormer [25], our method shows enhancements of 5.3 mm, 4.1 mm, and 0.7 mm in P1 error. Notably, in a series of actions, our method achieves the best performance recorded to date in 10 categories. When comparing with Motionbert [43] and MotionAGFormer [44] using the more powerful 2D detector Stacked Hourglass (SH) [46], our method further improves by approximately 2.3 mm (5.7%) in P1 error and 0.9 mm (2.8%) in P2 error. Our model significantly outperforms all other methods, achieving an approximate 0.4 mm improvement in average MPJPE compared to MotionAGFormer [44]. Table 3 shows the performance of our method on P1 error when using ground-truth 2D poses as input. Compared to CPN or SH 2D poses, the use of ground-truth 2D poses eliminates noise interference, enabling each model to reach peak performance. The trend of P1 error remains consistent, with our method leading all competitors in the 243-frame input case, achieving a P1 error of 21.4 mm, which outperforms the current leading model, MixSTE [20], by 0.2 mm.
Performance Comparison on MPI-INF-3DHP. To assess the generalizability of our proposed framework, we tested it on the MPI-INF-3DHP dataset, which is known for its challenging backgrounds. Given the shorter video sequences, we configured the number of input frames to be 9, 27, and 81, respectively. Detailed performance comparisons are provided in Table 4. As observed in the results on Human3.6M, the optimal performance is achieved with 81 frames, reaching a PCK of 98.8, AUC of 84.1, and P1 error of 19.2 mm. This performance outperforms the current state-of-the-art model by 0.1 in PCK and 0.2 in AUC. These results underscore the effectiveness of the proposed method in handling more complex scenarios. In our experiments, the MPI-INF-3DHP dataset primarily consists of complex outdoor scenes, with a focus on evaluating the accuracy of algorithm results. The experimental findings show that the accuracy metrics, AUC and PCK, outperform SOTA methods, including MotionAGFormer. On the other hand, for the precision metric MPJPE, comparisons on the Human3.6M dataset are more extensive, and our method’s MPJPE and P-MPJPE metrics outperform MotionAGFormer.
Table 1: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN and SH as inputs under Protocol 1. Methods marked with * denote transformer-based approaches, ^indicates GCN-based methods, and $\dagger$ represents graph-augmented transformer methods.
@@
StridedFormer [22]* (CPN, T=243) TMM’22 40.3 43.3 40.2 42.3 45.6 52.3 41.8 40.5 55.9 60.6 44.2 43.0 44.2 30.0 30.2 43.7
MHFormer [21]* (CPN, T=351) CVPR’22 39.2 43.1 40.1 40.9 44.9 51.2 40.6 41.3 53.5 60.3 43.7 41.1 43.8 29.8 30.6 43.0
MixSTE [20]* (CPN, T=243) CVPR’22 37.6 40.9 37.3 39.7 42.3 49.9 40.1 39.8 51.7 55.0 42.1 39.8 41.0 27.9 27.9 40.9
P-STMO [41]* (CPN, T=243) ECCV’22 38.9 42.7 40.4 41.1 45.6 49.7 40.9 39.9 55.5 59.4 44.9 42.2 42.7 29.4 29.4 42.8
HSTFormer [18]* (CPN, T=81) arXiv’23 39.5 42.0 39.9 40.8 44.4 50.9 40.9 41.3 54.7 58.8 43.6 40.7 43.4 30.1 30.4 42.7
Einfalt et al. [19]* (CPN, T=351) WACV’23 39.6 43.8 40.2 42.4 46.5 53.9 42.3 42.5 55.7 62.3 45.1 43.0 44.7 30.1 30.8 44.2
UGCN [42]^(CPN, T=96) ECCV’20 41.3 43.9 44.0 42.2 48.0 57.1 42.2 43.2 57.3 61.3 47.0 43.5 47.0 32.6 31.8 45.6
GLA-GCN [11]^(CPN, T=243) CVPR’23 41.3 44.3 40.8 41.8 45.9 54.1 42.1 41.5 57.8 62.9 45.0 42.8 45.9 29.4 29.9 44.4
STCFormer [25$\dagger$ (CPN, T=243) CVPR’23 39.6 41.6 37.4 38.8 43.1 51.1 39.1 39.7 51.4 57.4 41.8 38.5 40.7 27.1 28.6 41.0
STGFormer (Ours, CPN, T=81) 39.5 41.8 37.2 39.2 43.2 50.5 40.3 40.2 51.1 56.6 42.3 39.0 40.3 27.7 28.1 41.1
STGFormer (Ours, CPN, T=243) 38.4 40.7 37.1 38.1 42.7 49.9 39.0 39.3 50.4 56.0 41.4 38.3 39.1 26.9 28.1 40.3
Motionbert [43$\dagger$ (SH, T=243) ICCV’23 36.3 38.7 38.6 33.6 42.1 50.1 36.2 35.7 50.1 56.6 41.3 37.4 37.7 25.6 26.5 39.2
MotionAGFormer [44$\dagger$ (SH, T=243) WACV’24 36.4 38.4 36.8 32.9 40.9 48.5 36.6 34.6 51.7 52.8 41.0 36.4 36.5 26.7 27.0 38.4
STGFormer (Ours, SH, T=81) 38.5 39.1 36.2 37.8 42.1 49.9 36.4 35.7 50.8 54.5 41.1 37.1 38.0 26.8 26.1 39.3
STGFormer (Ours, SH, T=243) 35.9 37.6 35.4 36.6 40.5 48.0 35.7 34.2 49.7 52.4 39.6 36.8 37.2 24.9 25.7 38.0
Table 2: Performance Comparison on Human3.6M utilizes 2D poses estimated by CPN and SH as inputs under Protocol 2. Methods marked with * denote transformer-based approaches, ^indicates GCN-based methods, and $\dagger$ represents graph-augmented transformer methods.
@@
StridedFormer [22]* (CPN, T =243) TMM’22 32.7 35.5 32.5 35.4 35.9 41.6 33.0 31.9 45.1 50.1 36.3 33.5 35.1 23.9 25.0 35.2
MHFormer [21]* (CPN, T=351) CVPR’22 31.5 34.9 32.8 33.6 35.3 39.6 32.0 32.2 43.5 48.7 36.4 32.6 34.3 23.9 25.1 34.4
MixSTE [20]* (CPN, T=243) CVPR’22 30.8 33.1 30.3 31.8 33.1 39.1 31.1 30.5 42.5 44.5 34.0 30.8 32.7 22.1 22.9 32.6
P-STMO [41]* (CPN, T=243) ECCV’22 31.3 35.2 32.9 33.9 35.4 39.3 32.5 31.5 44.6 48.2 36.3 32.9 34.4 23.8 23.9 34.4
HSTFormer [18]* (CPN, T=81) arXiv’23 31.1 33.7 33.0 33.2 33.6 38.8 31.9 31.5 43.7 46.3 35.7 31.5 33.1 24.2 24.5 33.7
Einfalt et al. [19]* (CPN, T=351) WACV’23 32.7 36.1 33.4 36.0 36.1 42.0 33.3 33.1 45.4 50.7 37.0 34.1 35.9 24.4 25.4 35.7
UGCN [42]^(CPN, T=96) ECCV’20 32.9 35.2 35.6 34.4 36.4 42.7 31.2 32.5 45.6 50.2 37.3 32.8 36.3 26.0 23.9 35.5
GLA-GCN [11]^(CPN, T=243) CVPR’23 32.4 35.3 32.6 34.2 35.0 42.1 32.1 31.9 45.5 49.5 36.1 32.4 35.6 23.5 24.7 34.8
STCFormer [25$\dagger$ (CPN, T=243) CVPR’23 29.5 33.2 30.6 31.0 33.0 38.0 30.4 29.4 41.8 45.2 33.6 29.5 31.6 21.3 22.6 32.0
STGFormer (Ours, CPN, T=81) 29.6 33.2 30.2 31.3 33.0 37.7 30.1 30.4 41.7 44.6 33.5 29.5 31.7 22.0 23.1 32.1
STGFormer (Ours, CPN, T=243) 29.3 33.0 29.9 30.6 32.5 37.5 30.3 29.6 41.0 44.1 33.2 29.1 30.9 21.7 22.8 31.7
Motionbert [43$\dagger$ (SH, T=243) ICCV’23 30.8 32.8 32.4 28.7 34.3 38.9 30.1 30.0 42.5 49.7 36.0 30.8 22.0 31.7 23.0 32.9
MotionAGFormer [44$\dagger$ (SH, T=243) WACV’24 30.6 32.6 32.2 28.2 33.8 38.6 30.5 29.9 43.3 47.0 35.2 29.8 31.4 22.7 23.5 32.6
STGFormer (Ours, SH, T=81) 29.4 33.1 29.3 28.9 32.4 37.8 29.8 30.2 40.6 44.1 34.2 29.9 30.5 21.3 23.0 31.6
STGFormer (Ours, SH, T=243) 28.4 32.1 28.6 27.9 31.4 36.9 29.0 29.3 40.0 43.2 33.6 29.3 29.6 20.5 22.4 30.8
Table 3: Performance Comparison on Human3.6M uses ground-truth 2D poses as inputs under Protocol 1. Methods marked with * denote transformer-based approaches, ^indicates GCN-based methods, and $\dagger$ represents graph-augmented transformer methods.
@@
Table 4: Performance Comparison on MPI-INF-3DHP: Methods marked with * denote transformer-based approaches, ^indicates GCN-based methods, and $\dagger$ represents graph-augmented transformer methods.
@@
MotionAGFormer [44$\dagger$ (T=27) WACV’23 98.2 83.5 19.2
@@
In this section, we conduct a qualitative analysis of the model’s performance by visualizing the attention layers and assessing the effects of model on Human3.6M dataset and videos in-the-wild.
Attention visualization. The walking action from test set S11 is selected to demonstrate the attention mechanisms. The visualization of the attention layers shows the distribution of attention weights across different joints and time steps, with brighter colors indicating higher focus on the respective node during the current phase of motion. In the spatial attention map shown in Figure 4(a), both the horizontal and vertical axes represent joint indices. The dependencies between the left hip, knee, and ankle nodes and the right hip, knee, and ankle nodes are clearly significant for the "walking" sequence. The temporal attention layer in Figure 4(b) uses the x-axis and y-axis to represent different frames. This visualization reveals strong correlations between joints across adjacent frames as well as higher-order frames. Additionally, it shows that higher-order temporal connections, facilitated by self-attention, effectively capture joint temporal relationships from an alternative perspective.
Figure 4: Visualization of attention maps from the spatial and temporal attention modules in STGFormer.
Result visualization on Human3.6M. We also evaluate the visual results of the estimated poses and 3D ground truth from the Human3.6M dataset, comparing them with the state-of-the-art methods STCFormer [25] and MotionAGFormer [44] in Figure 5. The figure showcases four randomly selected examples from the Walking, Sitting, Greeting, and Eating sequences. Leveraging the framework’s robust modeling capabilities in both temporal and spatial dimensions, our method demonstrates strong performance in handling complex actions, such as Sitting (second row) and Eating (fourth row), by effectively reconstructing poses in three-dimensional space. These results highlight that our approach yields more accurate pose estimations compared to both STCFormer and MotionAGFormer.
@@
Result visualization on videos in-the-wild. Estimating 3D human poses from in-the-wild videos presents significant challenges due to complex environmental factors and unknown camera parameters. One practical way to assess a model’s generalization ability is by applying a pre-trained network to such videos. In this study, we first use YOLOv3 [47] to detect persons in the videos, followed by HRNet [48] for 2D keypoint detection. We then employ the pre-trained STGFormer model, trained on the Human3.6M dataset, to estimate the 3D human poses for videos sourced from Bilibili. For evaluation, we select a variety of challenging video clips, including fencing, dance, yoga, and figure skating. As shown in Figure 6, our method produces plausible and high-fidelity results for these in-the-wild videos, demonstrating the superior generalization capability of our approach.
Figure 6: Examples of 3D pose estimation by our framework on in-the-wild vidoes.
@@
To further investigate the performance contributions of different components of our proposed model, we conducted a series of ablation experiments on the Human3.6M dataset, using CPN-pretrained 2D keypoint coordinates as input. The ablation studies were organized into two main parts: (1) the impact of frame count on model performance, and (2) the effect of different model components on the overall accuracy.
Impact of frame count. In the first set of ablation studies, we explored how the performance of our proposed model varies with different frame counts. As illustrated in Table 5, the P1 error comparison of our model with other methods on the Human3.6M dataset is detailed for 27, 81, and 243 frames. Overall, we observed a trend of decreasing P1 error with an increasing frame count in our method and others. Among these methods, our model achieved the best results at 27, 81, and 243 frames, demonstrating strong performance in 3D pose estimation in long sequence videos. Notably, with a comparable number of parameters, our model requires less than 79% of the FLOPs compared to the currently best-performing MixSTE [20] at 243 frames.
Table 5: P1 error comparisons with different numbers of CPN-pretrained 2D keypoint frames on Human3.6M dataset. The best result in each column is highlighted in red.
@@
Analysis of model components. The second set of ablation studies examines the performance contribution of various components in our proposed model, specifically focusing on the effects of spatial and temporal graph attention, as well as different graph structures, as detailed in Table 6. In the table, STG attention (STGA) refers to the parallel combination of spatial graph attention (SGA) and temporal graph attention (TGA), which correspond to the spatial and temporal attention mechanisms in Section 3.3, respectively. SGA $\rightarrow$ TGA and TGA $\rightarrow$ SGA represent serial structures with reversed orders. MHR-GCN refers to the parallel dual-path structure consisting of S-GCN and T-GCN, as described in Section 3.4. S-GCN $\rightarrow$ T-GCN and T-GCN $\rightarrow$ S-GCN represent serial connections between the two components in different orders. GCN represents the case where no spatio-temporal splitting is performed.
Initially, when either SGA or TGA was applied independently with an MLP replacing the graph convolutional network, the P1 error increased significantly, showing values of 263.2 mm and 57.8 mm, respectively. This indicates that using attention mechanisms in isolation results in suboptimal feature representation, as both spatial and temporal dependencies need to be fully captured to improve the model’s performance. When combining the spatial and temporal graph attention (STGA) in parallel, the error was reduced to 53.3 mm, demonstrating the enhanced efficacy of jointly capturing both spatial and temporal dependencies. Further analysis reveals that the order of attention mechanisms also impacts the performance. The serial configuration of SGA $\rightarrow$ TGA resulted in a P1 error of 54.1 mm, while the reverse configuration, TGA $\rightarrow$ SGA, produced a higher error of 120.5 mm. This suggests that temporal dependencies are more effectively captured before spatial ones in our framework, emphasizing the importance of modeling temporal dynamics before refining spatial relationships. The introduction of the MHR-GCN, consisting of the spatial graph convolution network (S-GCN) and temporal graph convolution network (T-GCN) in parallel, leads to a further reduction in error to 43.7 mm. This result highlights the importance of both spatial and temporal graph convolutions working together in parallel to jointly capture complex dependencies, thereby providing more accurate feature representations. Interestingly, when combining S-GCN and T-GCN in serial (either S-GCN $\rightarrow$ T-GCN or T-GCN $\rightarrow$ S-GCN), the errors were slightly higher (52.5 mm and 47.0 mm, respectively), which suggests that parallel pathways are more effective than serial ones for processing spatial and temporal information simultaneously. These results indicate that each component plays a crucial role in improving model performance. The spatial and temporal graph attention mechanisms are essential for capturing the respective dependencies, while the parallel S-GCN and T-GCN structure offers a more robust representation of human pose dynamics compared to serial connections.
Table 6: Performance contribution of each component in the proposed framework on Human3.6M dataset.
attention part graph part P1 (mm)
SGA MLP 263.2
@@
SGA $\rightarrow$ TGA 54.1
TGA $\rightarrow$ SGA 120.5
STGA 53.3
STGA GCN 50.6
S-GCN 46.9
T-GCN 45.0
S-GCN $\rightarrow$ T-GCN 52.5
T-GCN $\rightarrow$ S-GCN 47.0
MHR-GCN 43.7
5 Conclusion
In this paper, we introduce Spatio-Temporal GraphFormer, a novel model that integrates long-range temporal-spatial dependencies with graph structural information in its attention mechanisms. It utilizes a dual-path modulated hop-wise regular GCN module to enhance graph information extraction for 3D human pose estimation in videos. Specifically, the spatio-temporal criss-cross graph attention separately models spatial and temporal correlations, leveraging encoded graph structure information. The dual-path modulated hop-wise regular GCN module adjusts graph weights and accounts for higher-order dependencies across both time and space dimensions. Experiments on two benchmark datasets affirm the efficacy and superior generalization capability of our proposed model compared to state-of-the-art techniques. While the proposed method presents several promising innovations for improving 3D human pose estimation, it still has limitations in multi-person and occlusion scenarios. First, the model assumes a single-person context, which could lead to performance degradation in frames with multiple individuals. The attention mechanism may struggle to effectively distinguish overlapping poses, leading to confusion in keypoint estimation. Additionally, occlusions, where parts of the body are hidden, present a significant challenge. Although spatiotemporal graph attention captures long-range dependencies, occlusions can disrupt the model’s ability to accurately infer 3D pose information, especially when joints are partially visible. Thus, while the method performs well in controlled, single-person scenarios, its robustness in crowded or occluded environments requires further exploration and enhancement. Addressing these limitations would improve the model’s applicability in diverse and practical settings. For future work, we aim to develop a multi-person skeletal sequence spatio-temporal graph representation method, extending our model’s applicability to 3D human pose estimation in multi-person scenarios and broadening its use across various contexts.